OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 2x − 3y = 3 and passes through the point (−8, 2).

OpenStudy (anonymous):

In this the slope of the given line is 2/3. So, the slope of the perpendicular line will be -3/2. So, line: y= -3x/2 + b When x = -8 then y=2: 2 = -3(-8)/2 + b => 2 = 12 + b=>b=-10. So, line is: y= -3x/2 - 10.